OpenStudy (anonymous):

Solve and check. Label any contradictions or identities. Please show all work. 5(x-7)=3(x-2)+2x

OpenStudy (accessdenied):

Do you know how to start the problem? I'm assuming this is 5*(x-7) = 3*(x-2) + 2x, correct me if I'm wrong...

OpenStudy (anonymous):

no i dont and you are right

OpenStudy (accessdenied):

Okay. Well, do you know distributive property? A(B+C) = AB + AC ? I can see this is applicable on this problem

OpenStudy (anonymous):

kinda

OpenStudy (accessdenied):

In this problem, we have 5(x - 7) and 3(x - 2). We can use distributive property here because we have a single quantity multiplied to two numbers inside parentheses. . If it helps to think about it this way, we can 'define' multiplication as 'adding a number some number of times.' Like.. 2(2+3) = (2+3) + (2+3) We have "2*...", so we add the thing we multiply twice = 2 + 3 + 2 + 3 = 10 2(2+3) = 2(5) If we simplify in the parentheses, we can see this is the same as the above = 10 So, back to that definition of distributing: 2(2+3) = 2*2 + 2*3 = 4 + 6 = 10 that's true! We can do the same thing if there's a variable in the thing we're multiplying

OpenStudy (anonymous):

how would I show that work in word

OpenStudy (accessdenied):

5(x-7) = (x-7) + (x-7) + (x-7) + (x-7) + (x-7) = 5x (adding 5 x's) - 35 (subtracting 5 7's) Distributive Property instead: 5(x - 7) = 5*x - 7*5 = 5x - 35 So, in this problem, we'd start by using distributive property... 5(x-7) = 3(x-2) + 2x 5x - 35 = 3x - 6 + 2x Distributive Property

OpenStudy (accessdenied):

like, explaining the process itself? or are you doing this on microsoft word..?

OpenStudy (anonymous):

Microsoft word

OpenStudy (accessdenied):

not sure... i guess if you do two column, you can just type out the equation on the left and then add the reasoning on the right..